For each of the unrelated transactions described below, present the entries required to record each transaction.

1. Martinez Corp. issued $19,500,000 par value 11% convertible bonds at 98. If the bonds had not been convertible, the company’s investment banker estimates they would have been sold at 95.
2. Sandhill Company issued $19,500,000 par value 11% bonds at 97. One detachable stock purchase warrant was issued with each $100 par value bond. At the time of issuance, the warrants were selling for $5.
3. Suppose Sepracor, Inc. called its convertible debt in 2020. Assume the following related to the transaction. The 12%, $10,500,000 par value bonds were converted into 1,050,000 shares of $1 par value common stock on July 1, 2020. On July 1, there was $57,000 of unamortized discount applicable to the bonds, and the company paid an additional $82,000 to the bondholders to induce conversion of all the bonds. The company records the conversion using the book value method.


(Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

No.

Account Titles and Explanation

Debit

Credit

1.

enter an account title for the first transaction

enter a debit amount

enter a credit amount

enter an account title for the first transaction

enter a debit amount

enter a credit amount

enter an account title for the first transaction

enter a debit amount

enter a credit amount

2.

enter an account title for the second transaction

enter a debit amount

enter a credit amount

enter an account title for the second transaction

enter a debit amount

enter a credit amount

enter an account title for the second transaction

enter a debit amount

enter a credit amount

enter an account title for the second transaction

enter a debit amount

enter a credit amount

3.

enter an account title for the third transaction

enter a debit amount

enter a credit amount

enter an account title for the third transaction

enter a debit amount

enter a credit amount

enter an account title for the third transaction

enter a debit amount

enter a credit amount

enter an account title for the third transaction

enter a debit amount

enter a credit amount

enter an account title for the third transaction

enter a debit amount

enter a credit amount

enter an account title for the third transaction

enter a debit amount

enter a credit amount
Answer #1

1)

Cash ($19,500,000 * 0.98) $19,110,000
Discount on Bonds Payable $390,000
Bonds Payable $19,500,000

2)

Cash $18,915,000
Discount on Bonds Payable $1,560,000
Bonds Payable $19,500,000
Paid-in Capital—Stock Warrants $975,000

Value of bonds plus warrents = 19,500,000 * 0.97 = 18,915,000

Value of warrants = 19,500,000 / 100 = 195,000 * $5 = $975,000

3)

Bonds Payable $10,500,000
Debt Conversion Expense $82,000
Discount on Bonds Payable $57,000
Common Stock $1,050,000
Paid-in Capital in Excess of Par ($10,500,000 - $57,000 - $1,050,000) $9,393,000
Cash $82,000